OpenStudy (anonymous):

PLEASE HELP ????Find the slope and determined by following 2 pair of points (2,-2), (1,2/3)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(2\quad ,&-2)\quad &(1\quad ,&\frac{2}{3}) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)